OpenStudy (anonymous):

find 3 irrational nubers between 4 &5

OpenStudy (solomonzelman):

\[4=\sqrt{16},~~~~~~~~5=\sqrt{25}\] 17, 18, 19, 20, 21, 22, 23, 24 \[example:~~~~~~~~\sqrt{18}\]

ganeshie8 (ganeshie8):

Clever :) another way is to use a known irrational number like \(\large \sqrt{2}\) : \(\large 4 + \dfrac{1}{k\sqrt{2}}\) \(k = 1,2,3....\) all will be infinite irrational numbers between 4 and 5 (wana prove ?)

OpenStudy (anonymous):

all irrational numbers are numbers that cannot be reduced into a simple fraction btw